OpenStudy (anonymous):

Write each logarithmic expression as a single logarithm. log 8 -2 log 6 + log 3 log 5 - k log 2

OpenStudy (anonymous):

\[\log 5 - k \log 2 = \log(5/2^{k})\]\[\log 8 -2 \log 6 + \log 3 =\log((8/36)*3)\]